# FlorDB: Multiversion Hindsight Logging for Continuous Training

### v3

@@ Abstract.
PVLDB Reference Format:
PVLDB, 17(1): XXX-XXX, 2024.
doi:XX.XX/XXX.XX This work is licensed under the Creative Commons BY-NC-ND 4.0 International License. Visit https://creativecommons.org/licenses/by-nc-nd/4.0/ to view a copy of this license. For any use beyond those covered by this license, obtain permission by emailing info@vldb.org. Copyright is held by the owner/author(s). Publication rights licensed to the VLDB Endowment.
Proceedings of the VLDB Endowment, Vol. 17, No. 1 ISSN 2150-8097.
doi:XX.XX/XXX.XX
PVLDB Artifact Availability:
The source code, data, and/or other artifacts have been made available at https://github.com/ucbrise/flor.
@@ 3.1. Running Experiments
Alice’s script (3) uses native Python functionality to run her ML experiment. Hyper-parameters such as learning rate and batch size are passed in using standard command-line arguments, which makes hindsight logging seamless and capable of interoperating with other tools typically used in the ML workflow.
@@ 3.1.1. Args & Logs
In the context of running experiments, Alice utilizes the Flor API to enhance configurability and traceability. The flor.arg function (lines 5-8 in 3) is instrumental in parameterizing her experiments, allowing her to define hyper-parameters such as hidden_size, learning_rate, batch_size, and seed either via the command-line or by utilizing default values specified within the script. This ensures adaptability and reproducibility of her experiment’s configuration. In addition, flor.log is employed (line 20 in 3) to record the experiment’s loss metrics and associated metadata at each optimizer step. The logging of the loss captures the dynamics of the model’s performance, enabling Alice to monitor the training process.
@@ 3.2. Experiment Analysis and Hindsight Logging
Alice reviews FlorDB’s flor.dataframe output, comparing F1-scores across datasets (see 4). After receiving insights from her colleague, Alice shifts her analysis focus to the pedestrian F1-scores. By editing train.py to include a new flor.log for f1_ped, Alice leverages hindsight logging to populate this metric retroactively. She invokes the replay feature via CLI as follows:
@@
Upon initiating the replay command, Alice is shown a schedule of experiments slated for replay and must confirm to proceed. After receiving confirmation, Flor prepares to replay the selected prior runs: it restores their versions from github, adds the new f1_ped logging statement to each, and executes all versions using the configuration variable values originally logged via flor.arg. Notably, during this replay, Flor’s hindsight logging functionality can bypass the main loop (lines 12-22 in 3) to directly load the model’s final state, since the f1_ped statement is computed post-training (line 25 in 3).
@@
This command now generates a table including the f1_ped metric (green text in 4) for entries dated June 23, 2023, or later. This new data reveals a previously unnoticed performance regression in pedestrian classifications, as Bob had indicated (f1_ped values in the range tstamp $\geq$ 08-27).
@@ 4. System Architecture
The design of FlorDB aims to provide a unified platform for the management of code, checkpoints, logs, and their evolution over time (5). FlorDB maintains a cross-referenced relational database that interconnects code versions in git, large files in your preferred storage solution (e.g. an S3 bucket), and metadata stored in the database. This section outlines the mechanisms FlorDB uses for executing experiments and replaying them for hindsight logging, highlighting how it stores, manages and tracks the data and metadata of model training.
@@ 4.1. Storage and Data Layout
As shown in the bottom of 5, FlorDB’s storage architecture consists of the following three units:
@@ 4.2. Relational Schema
The relational schema of FlorDB is a structured and normalized view over the JSON logs. As depicted in 6, the logs table is central to this schema, managing the data and relevant metadata of experiments. The schema for the logs table includes:
@@
pt
ctx_id: integer parent_ctx_id: integer loop_name: text loop_entries: integer loop_iteration: integer
pt
@@
pt
projid: text tstamp: datetime filename: text ctx_id: integer value_name: text value: text value_type: integer
pt
@@
pt
ts_start: datetime vid: text ts_end: datetime
pt
@@
pt
vid: text filename: text parent_vid: text contents: text
pt
@@
pt
projid: text tstamp: datetime filename: text ctx_id: integer value_name: text contents: blob
@@ 4.3. Pivoted Views
The flor.dataframe presents the content of the database in a single-table view (7). This view is generated from the relational schema (6) by joining the logs table with additional dimensions of data, such as the loops table, which are essential for capturing the full context of each experimental run. This is followed by a series of transformations that pivot the combined data, turning relational values (e.g., logs.value_name, loops.loop_name) into column headers, to create a wide-format table where each unique loop and log entry name becomes a distinct column in the DataFrame (e.g. 4).
@@ 4.3.1. Domain Mapping
  • Running Experiments normal-$\rightarrow$ Adding Rows: The execution of a new experiment results in the addition of rows to flor.dataframe. Each row in the DataFrame signifies a discrete iteration within an experiment, collating related data such as metrics, hyper-parameters, and state.
  • Adding Logging Statements normal-$\rightarrow$ Adding Virtual Columns: The inclusion of logging statements in the source code induces FlorDB to add virtual columns to the view produced by flor.dataframe. These virtual columns can then be referenced in subsequent queries like any other column in flor.dataframe.
  • Replay from Checkpoint normal-$\rightarrow$ Backfilling Nulls: Replaying an experiment from checkpoints corresponds to the backfilling of null values in the flor.dataframe view. This capability is critical to achieving the hindsight logging abstraction that FlorDB supports.
@@ 5. Acquisitional Query Processing
FlorDB extends the concept of Acquisitional Query Processing (AQP) (Madden et al., 2005), which involves efficient data acquisition during query execution. AQP was originally presented in the context of live sensing; here we acquire data by re-running training code from checkpoints. The process begins at the query construction stage, where criteria for selecting experimental versions are established (5.1). Each selected version has logging statements propagated to it in order to generate the required data (5.3). Subsequently, a modified training script is executed with a --replay_flor flag, enabling replay query operators to perform partial or parallel model training replay (5.2.1). In this section, we describe multiversion hindsight logging via the sequence of mechanisms by which FlorDB integrates data acquisition into query execution. This also paves the way for comprehensive post-hoc analysis and continuous model refinement.
@@ 5.2. Empirical Cost Prediction & Plan Execution
Machine Learning (ML) training scripts can be long-running, and even selective replay can be time-consuming. Therefore, ML engineers need to know how long a replay will take, to decide whether further query refinement is needed. Traditional query cost estimation is a notoriously hard problem, often exhibiting orders of magnitude errors in estimation (Leis et al., 2015). By contrast, we can estimate query runtime for replay executions extremely accurately (3). This is because FlorDB profiles runtime information during the record phase. When initiating a Flor replay, these runtime statistics are queried using the flor.dataframe API call. This allows us to reliably predict the time required for replay, enabling users to tailor their queries for efficiency. We categorize the cost estimation into four levels:
@@
    Prefix Scan: Executes statements before the main loop, logging preliminary setup operations (e.g., lines 1-10 in 3. This level is useful when examining data preparation and featurization.
@@
    Suffix Scan: Executes the initial setup code (as in prefix scan), loads the end state of the outermost loop from a checkpoint, and runs the final script segments (lines 23-EOF in 3). This level provides a more detailed view, useful when examining the results of training (e.g. accuracy, recall, F1-scores).
@@
    Validation Scan. Steps through the main training loop to execute model validation logic (lines 21-22 in 3)., loading checkpoints once per epoch but skipping the nested training loop. This level is useful for examining the validation process in detail.
@@
    Range Scan: Runs the training loop in depth (lines 12-22 in 3) over a selected range to capture nested logging statements — used for the most granular analyses (e.g. of gradients). A range scan running from epoch 0 to $N$ is considered a full scan. To achieve replay parallelism, FlorDB executes range scans over non-overlapping intervals.
@@ 5.3. Logging Statement Propagation
FlorDB’s code alignment uses GumTree (Falleri et al., 2014) for aligning code blocks, aiding in the accurate insertion of log statements across versions. The flor.loop() function establishes stable anchor points, with main and nested training loops typically serving as consistent markers across versions. This consistency, enhanced by GumTree’s alignment capabilities, ensures accurate log statement propagation. While GumTree is capable of aligning code blocks without Flor API anchors, the presence of these static loop identifiers all but ensures its efficiency and accuracy: in all instances of logging statement propagation evaluated in 6, this combined approach was successful.
@@ 6.1. Efficient Linear Scaling across Versions
To assess FlorDB’s ability to manage multiversion hindsight logging without cross-version interference, we evaluated its performance across different model architectures. For each model architecture, we measured the runtime for the following modes of execution (as defined in 5.2):
@@
Our findings, depicted in 9, reveal that FlorDB’s time complexity scales linearly with an increasing number of versions, as demonstrated by the log-scale on the y-axis. This evidence of linear scalability, devoid of interaction effects or incremental overhead, validates FlorDB’s ability to handle both large models and big datasets.
@@ 6.1.1. Interactive Response Times
In evaluating the impact of different replay operations on system response times (see orange lines in 9), the replay-prefix operation maintains the fastest processing times, consistently delivering results in under 10 seconds. Even as the number of versions increases, the processing time for replay-prefix remains the most stable, ensuring a responsive user experience. In contrast, replay-suffix and replay-validation operations exhibit longer processing times as the version count rises, with replay-validation times increasing more steeply. One factor is the amount of time it takes to evaluate the model. While replay-suffix may still offer interactive speeds at lower version counts, replay-validation quickly surpasses the interactive threshold: in scenarios where immediate feedback is crucial, users can refine the selectivity of their queries for faster response times (as discussed in 5.2).
@@ 6.1.2. Storage Requirements
Despite frequent commits across numerous versions, Git repositories remain remarkably compact, typically under 5MB (see right pane of 10). In contrast, checkpoints for large models stored on shared storage can easily consume hundreds of gigabytes (see orange bars in the left pane). To address potential local storage depletion, FlorDB’s default approach is to spool least-recently-used checkpoints to a cloud object store like S3444Storing 100 GB of data in S3 costs approximately \$2.30 a month.. If this is not feasible, FlorDB may retain a subset of checkpoint per version—as few as just two (see green bars in 10)—evicting the rest. FlorDB retains checkpoints that are evenly spaced across iterations. This ensures balanced work partitions and avoids stragglers that would hinder replay parallelism. Importantly, even retaining just two checkpoints (one mid-run, one at completion) offers significant benefits: i) the final checkpoint enables fast restores for replay-prefix and replay-suffix queries, and the mid-run checkpoint enables 2x faster replay execution compared to the original run.
@@ 6.2.1. Bottleneck Analysis — length-1 range scan
11 (top) depicts the runtime of a range scan over a single epoch. Assuming unbounded resources (e.g. in an idealized cloud), every version would be replayed in its own machine. In this ideal scenario, a query’s bottleneck would be the time it takes for a single Flor replay operation to finish. Replay-prefix, replay-suffix, and replay-validation would each finish in about 100 seconds or less (9). In contrast, the cost of a range scan, even for a single epoch, is higher. This operation requires stepping into the nested training loop, performing a forward pass over the neural network, and back-propagating gradients. Consequently, the range scan emerges as a bottleneck operator. In 11 we evaluate its runtime on a single epoch; running times for more epochs or versions can be linearly extrapolated. This is due to the ideal parallelism of replay-from-checkpoint (confirmed by prior studies (Garcia et al., 2020)) and the embarrassing parallelism of replay across versions.
@@ 6.2.2. Cost and Performance Trade-offs
Our bottleneck analysis assumed an ideal of unbounded resources. While parallelism can significantly accelerate computation, it comes with increased resource consumption and financial costs. 11 (bottom) compares the estimated cost per epoch across CPU and GPU platforms. Costs are based on AWS EC2 instance rental fees, with and without GPUs, factoring in execution time. GPUs provide order-of-magnitude superior computational speed but incur order-of-magnitude higher expenses. CPUs have the added benefit that they can be elastically allocated en masse. This can improve throughput across many versions or epochs, but as the bottleneck analysis indicates, there are limits to response times when range scans are used.

### v4

# Multiversion Hindsight Logging for Continuous Training

@@ 3.1. Running Experiments
Alice’s script (Figure 3) uses native Python functionality to run her ML experiment. Hyper-parameters such as learning rate and batch size are passed in using standard command-line arguments, which makes hindsight logging seamless and capable of interoperating with other tools typically used in the ML workflow.
@@ 3.1.1. Args & Logs
In the context of running experiments, Alice utilizes the Flor API to enhance configurability and traceability. The flor.arg function (lines 5-8 in Figure 3) is instrumental in parameterizing her experiments, allowing her to define hyper-parameters such as hidden_size, learning_rate, batch_size, and seed either via the command-line or by utilizing default values specified within the script. This ensures adaptability and reproducibility of her experiment’s configuration. In addition, flor.log is employed (line 20 in Figure 3) to record the experiment’s loss metrics and associated metadata at each optimizer step. The logging of the loss captures the dynamics of the model’s performance, enabling Alice to monitor the training process.
@@ 3.2. Experiment Analysis and Hindsight Logging
Alice reviews FlorDB’s flor.dataframe output, comparing F1-scores across datasets (see Figure 4). After receiving insights from her colleague, Alice shifts her analysis focus to the pedestrian F1-scores. By editing train.py to include a new flor.log for f1_ped, Alice leverages hindsight logging to populate this metric retroactively. She invokes the replay feature via CLI as follows:
@@
Upon initiating the replay command, Alice is shown a schedule of experiments slated for replay and must confirm to proceed. After receiving confirmation, Flor prepares to replay the selected prior runs: it restores their versions from github, adds the new f1_ped logging statement to each, and executes all versions using the configuration variable values originally logged via flor.arg. Notably, during this replay, Flor’s hindsight logging functionality can bypass the main loop (lines 12-22 in Figure 3) to directly load the model’s final state, since the f1_ped statement is computed post-training (line 25 in Figure 3).
@@
This command now generates a table including the f1_ped metric (green text in Figure 4) for entries dated June 23, 2023, or later. This new data reveals a previously unnoticed performance regression in pedestrian classifications, as Bob had indicated (f1_ped values in the range tstamp $\geq$ 08-27).
@@ 4. System Architecture
The design of FlorDB aims to provide a unified platform for the management of code, checkpoints, logs, and their evolution over time (Figure 5). FlorDB maintains a cross-referenced relational database that interconnects code versions in git, large files in your preferred storage solution (e.g. an S3 bucket), and metadata stored in the database. This section outlines the mechanisms FlorDB uses for executing experiments and replaying them for hindsight logging, highlighting how it stores, manages and tracks the data and metadata of model training.
@@ 4.1. Storage and Data Layout
As shown in the bottom of Figure 5, FlorDB’s storage architecture consists of the following three units:
@@ 4.2. Relational Schema
The relational schema of FlorDB is a structured and normalized view over the JSON logs. As depicted in Figure 6, the logs table is central to this schema, managing the data and relevant metadata of experiments. The schema for the logs table includes:
@@
ctx_id: integer
parent_ctx_id: integer
loop_name: text
loop_entries: integer
loop_iteration: integer
@@
projid: text
tstamp: datetime
filename: text
ctx_id: integer
value_name: text
value: text
value_type: integer
@@
ts_start: datetime
vid: text
ts_end: datetime
@@
vid: text
filename: text
parent_vid: text
contents: text
@@
projid: text
tstamp: datetime
filename: text
ctx_id: integer
value_name: text
contents: blob
@@ 4.3. Pivoted Views
The flor.dataframe presents the content of the database in a single-table view (Figure 7). This view is generated from the relational schema (Figure 6) by joining the logs table with additional dimensions of data, such as the loops table, which are essential for capturing the full context of each experimental run. This is followed by a series of transformations that pivot the combined data, turning relational values (e.g., logs.value_name, loops.loop_name) into column headers, to create a wide-format table where each unique loop and log entry name becomes a distinct column in the DataFrame (e.g. Figure 4).
@@ 4.3.1. Domain Mapping
  • Running Experiments $\rightarrow$ Adding Rows: The execution of a new experiment results in the addition of rows to flor.dataframe. Each row in the DataFrame signifies a discrete iteration within an experiment, collating related data such as metrics, hyper-parameters, and state.
  • Adding Logging Statements $\rightarrow$ Adding Virtual Columns: The inclusion of logging statements in the source code induces FlorDB to add virtual columns to the view produced by flor.dataframe. These virtual columns can then be referenced in subsequent queries like any other column in flor.dataframe.
  • Replay from Checkpoint $\rightarrow$ Backfilling Nulls: Replaying an experiment from checkpoints corresponds to the backfilling of null values in the flor.dataframe view. This capability is critical to achieving the hindsight logging abstraction that FlorDB supports.
@@ 5. Acquisitional Query Processing
FlorDB extends the concept of Acquisitional Query Processing (AQP) (Madden et al., 2005), which involves efficient data acquisition during query execution. AQP was originally presented in the context of live sensing; here we acquire data by re-running training code from checkpoints. The process begins at the query construction stage, where criteria for selecting experimental versions are established (Section 5.1). Each selected version has logging statements propagated to it in order to generate the required data (Section 5.3). Subsequently, a modified training script is executed with a --replay_flor flag, enabling replay query operators to perform partial or parallel model training replay (Section 5.2.1). In this section, we describe multiversion hindsight logging via the sequence of mechanisms by which FlorDB integrates data acquisition into query execution. This also paves the way for comprehensive post-hoc analysis and continuous model refinement.
@@ 5.2. Empirical Cost Prediction & Plan Execution
Machine Learning (ML) training scripts can be long-running, and even selective replay can be time-consuming. Therefore, ML engineers need to know how long a replay will take, to decide whether further query refinement is needed. Traditional query cost estimation is a notoriously hard problem, often exhibiting orders of magnitude errors in estimation (Leis et al., 2015). By contrast, we can estimate query runtime for replay executions extremely accurately (Footnote 3). This is because FlorDB profiles runtime information during the record phase. When initiating a Flor replay, these runtime statistics are queried using the flor.dataframe API call. This allows us to reliably predict the time required for replay, enabling users to tailor their queries for efficiency. We categorize the cost estimation into four levels:
@@
    Prefix Scan: Executes statements before the main loop, logging preliminary setup operations (e.g., lines 1-10 in Figure 3. This level is useful when examining data preparation and featurization.
@@
    Suffix Scan: Executes the initial setup code (as in prefix scan), loads the end state of the outermost loop from a checkpoint, and runs the final script segments (lines 23-EOF in Figure 3). This level provides a more detailed view, useful when examining the results of training (e.g. accuracy, recall, F1-scores).
@@
    Validation Scan. Steps through the main training loop to execute model validation logic (lines 21-22 in Figure 3)., loading checkpoints once per epoch but skipping the nested training loop. This level is useful for examining the validation process in detail.
@@
    Range Scan: Runs the training loop in depth (lines 12-22 in Figure 3) over a selected range to capture nested logging statements — used for the most granular analyses (e.g. of gradients). A range scan running from epoch 0 to $N$ is considered a full scan. To achieve replay parallelism, FlorDB executes range scans over non-overlapping intervals.
@@ 5.3. Logging Statement Propagation
FlorDB’s code alignment uses GumTree (Falleri et al., 2014) for aligning code blocks, aiding in the accurate insertion of log statements across versions. The flor.loop() function establishes stable anchor points, with main and nested training loops typically serving as consistent markers across versions. This consistency, enhanced by GumTree’s alignment capabilities, ensures accurate log statement propagation. While GumTree is capable of aligning code blocks without Flor API anchors, the presence of these static loop identifiers all but ensures its efficiency and accuracy: in all instances of logging statement propagation evaluated in Section 6, this combined approach was successful.
@@ 6.1. Efficient Linear Scaling across Versions
To assess FlorDB’s ability to manage multiversion hindsight logging without cross-version interference, we evaluated its performance across different model architectures. For each model architecture, we measured the runtime for the following modes of execution (as defined in Section 5.2):
@@
Our findings, depicted in Figure 9, reveal that FlorDB’s time complexity scales linearly with an increasing number of versions, as demonstrated by the log-scale on the y-axis. This evidence of linear scalability, devoid of interaction effects or incremental overhead, validates FlorDB’s ability to handle both large models and big datasets.
@@ 6.1.1. Interactive Response Times
In evaluating the impact of different replay operations on system response times (see orange lines in Figure 9), the replay-prefix operation maintains the fastest processing times, consistently delivering results in under 10 seconds. Even as the number of versions increases, the processing time for replay-prefix remains the most stable, ensuring a responsive user experience. In contrast, replay-suffix and replay-validation operations exhibit longer processing times as the version count rises, with replay-validation times increasing more steeply. One factor is the amount of time it takes to evaluate the model. While replay-suffix may still offer interactive speeds at lower version counts, replay-validation quickly surpasses the interactive threshold: in scenarios where immediate feedback is crucial, users can refine the selectivity of their queries for faster response times (as discussed in Section 5.2).
@@ 6.1.2. Storage Requirements
Despite frequent commits across numerous versions, Git repositories remain remarkably compact, typically under 5MB (see right pane of Figure 10). In contrast, checkpoints for large models stored on shared storage can easily consume hundreds of gigabytes (see orange bars in the left pane). To address potential local storage depletion, FlorDB’s default approach is to spool least-recently-used checkpoints to a cloud object store like S3444Storing 100 GB of data in S3 costs approximately \$2.30 a month.. If this is not feasible, FlorDB may retain a subset of checkpoint per version—as few as just two (see green bars in Figure 10)—evicting the rest. FlorDB retains checkpoints that are evenly spaced across iterations. This ensures balanced work partitions and avoids stragglers that would hinder replay parallelism. Importantly, even retaining just two checkpoints (one mid-run, one at completion) offers significant benefits: i) the final checkpoint enables fast restores for replay-prefix and replay-suffix queries, and the mid-run checkpoint enables 2x faster replay execution compared to the original run.
@@ 6.2.1. Bottleneck Analysis — length-1 range scan
Figure 11 (top) depicts the runtime of a range scan over a single epoch. Assuming unbounded resources (e.g. in an idealized cloud), every version would be replayed in its own machine. In this ideal scenario, a query’s bottleneck would be the time it takes for a single Flor replay operation to finish. Replay-prefix, replay-suffix, and replay-validation would each finish in about 100 seconds or less (Figure 9). In contrast, the cost of a range scan, even for a single epoch, is higher. This operation requires stepping into the nested training loop, performing a forward pass over the neural network, and back-propagating gradients. Consequently, the range scan emerges as a bottleneck operator. In Figure 11 we evaluate its runtime on a single epoch; running times for more epochs or versions can be linearly extrapolated. This is due to the ideal parallelism of replay-from-checkpoint (confirmed by prior studies (Garcia et al., 2020)) and the embarrassing parallelism of replay across versions.
@@ 6.2.2. Cost and Performance Trade-offs
Our bottleneck analysis assumed an ideal of unbounded resources. While parallelism can significantly accelerate computation, it comes with increased resource consumption and financial costs. Figure 11 (bottom) compares the estimated cost per epoch across CPU and GPU platforms. Costs are based on AWS EC2 instance rental fees, with and without GPUs, factoring in execution time. GPUs provide order-of-magnitude superior computational speed but incur order-of-magnitude higher expenses. CPUs have the added benefit that they can be elastically allocated en masse. This can improve throughput across many versions or epochs, but as the bottleneck analysis indicates, there are limits to response times when range scans are used.